In the laboratory, a student dilutes 11.0 mL of a 7.94 M hydrobromic acid solution to a total volume of 300.0 mL. What is the concentration of the diluted solution?

How many milliliters of 9.56 M hydrochloric acid solution should be used to prepare 2.50 L of 0.500 M HCL?

In the laboratory, a student adds 56.5 mL of water to 13.2 mL of a 0.875 M nitric acid solution. What is the concentration of the diluted solution? Assume the volumes are additive.
Answer #1

1)

use dilution formula

M1*V1 = M2*V2

1---> is for stock solution

2---> is for diluted solution

Given:

M1 = 7.94 M

V1 = 11 mL

V2 = 300 mL

use:

M1*V1 = M2*V2

M2 = (M1*V1)/V2

M2 = (7.94*11)/300

M2 = 0.2911 M

Answer: 0.291 M

2)

use dilution formula

M1*V1 = M2*V2

1---> is for stock solution

2---> is for diluted solution

Given:

M1 = 9.56 M

M2 = 0.5 M

V2 = 2500 mL

use:

M1*V1 = M2*V2

V1 = (M2 * V2) / M1

V1 = (0.5*2500)/9.56

V1 = 131 mL

Answer: 131 mL

3)

use dilution formula

M1*V1 = M2*V2

1---> is for stock solution

2---> is for diluted solution

Given:

M1 = 0.874 M

V1 = 13.2 mL

V2 = 13.2 mL + 56.5 mL = 69.7 mL

use:

M1*V1 = M2*V2

M2 = (M1*V1)/V2

M2 = (0.874*13.2)/69.7

M2 = 0.1655 M

Answer: 0.166 M